Project

General

Profile

8/12/14

Attendance

Matthew, Jonathan, Joao, Craig, Chris, Susan, Nick, Dominick, Nate, Michael, Jeny, Satish, Kanika

News (Matthew)

  • Production testing now runs nightly development tests (have all failed so far!). Will be expanded to first-ana development this week, then the S/FA tags when the next tag is available. They also now automatically archive the results after a certain amount of time.
  • IFDH spill info - beam quality cuts have been updated. These will be included in the next FA tag.
  • We may tag later this week after the current development issues are resolved.
  • Metrics and cookbook websites changed location, see email.

FD data (Dominick, Chris)

FD cosmics are still in LEM (with about a week to go), FD numi are ready for LEM when the cosmics are done. Chris will test how long the numi files will take. There was some discussion about the whether preselection is still being brought up at nu_e to speed future processing. It was noted that it wasn't. Matthew will communicate with Dominick about some inconsistencies observed in the FD cosmic data datasets.

ND data (Dominick)

The post-shutdown files are through PIDpart and LEM has been run. Although it looks like LEM was run in the wrong release maybe? Chris and Dominick will communicate about this.

Before we can make MC Barnali needs to make the list of files. Matthew will communicate this to DQ and to Ryan.

FD cry file replacement (Nate)

The artdaq processing is complete. It was noted that the number of pclist files doesn't correspond to either the number of files in the new dataset or the number in the old dataset. Nate will check the origin of these files and retire them if needed. Once this is complete Paola can restart her pclist cron jobs. The artdaq files are ready for reco and Matthew will pass the dataset name on to Satish.

There was some discussion about whether we should revisit the need for a uniquifier. Matthew will follow up on this.

FD 14DB genie RHC reco (Satish)

All done, except a handful of non-swap files that failed. Satish will take care of these today. Nate will retire the old daq files which will trim out the duplicate files.

Nate was also requested to put his retiring script into SVN.

Reco keep up (Jeny)

This was launched last week, and although there were some troubles at the beginning with the ND, things are now working smoothly (except some new troubles with with DAQ DB). One month worth of post-shutdown ND data has been processed as have one month worth of FD data. The FD data will be sent to Bruno to validate and the ND set will be topped up to the entire available dataset. Once these files have been signed off the cron job can be started.

DQ race condition and fixes (Joao et al.)

A script has been put in place which should fix files for which this is a problem. When it was run for the first time Joao found 24k files where this problem was present!He fixed most of the 24k files. But 107 files still fail. Adam may have figured out why this happened.

Renaming DAQ files (Jeny)

This has been done, and it shouldn't need to be done again.

Log file archival (Paola)

Paola is looking at this now and will update us as she finds out more.

AOB (all)

  • Any news on joining DBs to SAM? - Matthew will poke that thread.